OpenStudy (anonymous):

Find the exact value of tan(23degrees)+tan(7degrees)/1-tan(23degrees)tan(7degrees)

Parth (parthkohli):

seems like an application of the formula\[\tan(\alpha + \beta) = \dfrac{\tan \alpha + \tan\beta}{1 - \tan\alpha \tan \beta}\]